OpenStudy (vuriffy):

It is given that ln(y+1) - lny = 1 +3lnx. Express y in terms of x, in a form not involving logarithms.

OpenStudy (vuriffy):

@3mar

OpenStudy (vuriffy):

What I have so far is: ln(y+1) - lny = lne + 3lnx

OpenStudy (3mar):

Can you try firstly yourself? I am with you if you stuck!

OpenStudy (vuriffy):

Is it possible to cancel out all of the lns? Or am I just thinking very off? >.<

OpenStudy (vuriffy):

Oh, that wouldn't work, because it would cancel out the y. Yikes.

OpenStudy (vuriffy):

Should I do something with ln(y+1)?

OpenStudy (vuriffy):

Like distribute it?

OpenStudy (3mar):

You can't as there are operations between them Only you can if there are like that\[\ln(ffremlkfergknrel)=\ln(fewferfiop$Y%$)\] then \[ffremlkfergknrel=fewferfiop$Y%$\]

OpenStudy (vuriffy):

We are trying to isolate y, correct?

OpenStudy (vuriffy):

So, I would e both sides?

OpenStudy (3mar):

No, you would firstly do that: \[\ln(y+1) - lny = 1 +3lnx.\] \[\ln[\frac{ y+1 }{ y }]= 1 +3lnx\] agree?

OpenStudy (vuriffy):

OH because subtraction is division.

OpenStudy (3mar):

Excellent!

OpenStudy (3mar):

Then it would better to do that \[\ln[\frac{ y+1 }{ y }]= 1 +3lnx\] then \[\ln[\frac{ y+1 }{ y }]-3lnx=1\] then \[\ln[\frac{ y+1 }{ y }]-lnx^3=1\] then \[\ln[\frac{ y+1 }{ y*x^3 }]=1\] agree?

OpenStudy (vuriffy):

Well, I just wrote down the 3rd thing you had so I guess I agree. :o

OpenStudy (3mar):

Great! Now you can eliminate the ln by doing that \[\ln[\frac{ y+1 }{ y*x^3 }]=1\] then \[\ln[\frac{ y+1 }{ y*x^3 }]=\ln(e)\] then \[\frac{ y+1 }{ y*x^3 }=e\] then \[ y+1=e* y*x^3\] then \[ y+1=e*x^3*y\] then \[ y-e*x^3*y=-1\] then \[ y(1-e*x^3)=-1\] then \[\Huge y=\frac{ -1 }{ (1-e*x^3) }\]

OpenStudy (3mar):

Do you have any difficulties?

OpenStudy (vuriffy):

Let me check my work. I got that answer though.

OpenStudy (vuriffy):

Alright, all my steps were closely the same as yours so I got that too. That's it for it?

OpenStudy (3mar):

Do you need any more simplicity than this? ;]

OpenStudy (3mar):

@Vuriffy Satisfied?

OpenStudy (vuriffy):

A little confused, http://www.wolframalpha.com/input/?i=ln(y%2B1)+-+lny+%3D+1+%2B3lnx says the answer would be: y = 1/(ex^3 - 1)

OpenStudy (vuriffy):

Should the y be negative at any point, because if it was then you would multiply each side by -1 and it would equal that.

OpenStudy (vuriffy):

Ahh, they equal the same don't they?

OpenStudy (3mar):

\[\Large y=\frac{ -1 }{ (1-e*x^3) }=\frac{ +1 }{ (-1+e*x^3) }=\frac{ +1 }{ (e*x^3-1) }\] HAPPY NOW?

OpenStudy (vuriffy):

Haha, yes. Thank you so much. (:

OpenStudy (3mar):

Thank you for learning! Good learner! I really like who ask and stop the explanation when he sticks to something, even it looks so simple! Any more questions?

OpenStudy (vuriffy):

No, I am fine now! I appreciate it.